6w-(-a)=4+9w
2w-(-7a)=8+9a
W=
A=

Answer by ptth176(50)

6w-(-a)=4+9w=>a=3w+4(1)
2w-(-7a)=8+9a=>2w=8+2a(2)
(1)(2)=>2w=8+2(3w+4)
=>2w=8+6w+8
=>-4w=16
=>w=-4
(1)=>a=3(-4)+4=-8